Find the inverse Laplace transform
a . 3 s4 - 2s s2 (3s2 + 4) 3 S4 – 25 (s + 1)(3s2 + 4s + 2) b.
Answer #1

S (33 - 2) 3st – 25 S21352 +4) S2 (32+4) ir not les Since the degree of the denominator than degree of the denominator of per45 - 2 3 2. 3 3 s (BS+C)+ (35244 ) A 2 Ass 52 ( 3A+387 +350 +48 3 euating the coefficient 3A+ 3B=0 and A=-B 36=-2/2 and 4A= -- 4S- let F(5)= 3 s(52+ 4/5 35²44 7 3876 - 5 / 5 s 4 3 Ś Š + 1/2 32 44/12 FCS) 52462 2/35 52423 1-4*++ inga sistema both side